find the equation of the straight line passing through ((2/5),(3/7)) and perpendicular to y+8x=2x+7
Answer by Alan3354(69443)

find the equation of the straight line passing through ((2/5),(3/7)) and perpendicular to y+8x=2x+7
----------------
1st, find the slope of y+8x=2x+7.
y = -6x+7
That's the "slope intercept" form, y = mx+b, so the slope is -6.
The slope of a line perpendicular will have slope of the negative inverse, or 1/6.
Use that and the point to find the equation.
y-y1 = m*(x-x1)
y-3/7 = (1/6)*(x - 2/5)
Multiply by 210 to get rid of fractions
210y - 90 = 35*(x - 2/5)
210y - 90 = 35x -14
35x - 210y = -76 (standard form)
y = (1/6)x + 38/105 (slope intercept form)